the equation Y =4 sin pi(200t - X / 25) represents a transfers wave that travels in a streched wire x y is in cm and T in seconds its wavelength and velocity are

Grade:12th pass

1 Answers

Khimraj
5 years ago
k = 1/25
2\pi/\lambda = 1/25
\lambda = 50\pi cm
w = 200
v = w/k
v = 200*25 = 5000cm/s = 50m/s.
Hope it clears......................